{F} LA. find the minimal polynomial of the linear operator t : R³ "- R³" define by t (x,y,z) =(x+2y+3z, 4y+5z,6 z).is t


1
Expert's answer
2022-02-02T08:09:19-0500

Transforming matrix "t" is given by




"t=\begin{pmatrix}
 1&2&3 \\
 0&4&5 \\
0&0&6
\end{pmatrix}"


Characteristic polynomial of "t" is of form


"\lambda^3-D_1\lambda^2+D_2\lambda-D_3=0"



Where "D_1=" Sum of main diagonal element

"=1+4+6=11"


"D_2=" Sum of minors of the main diagonal element


"D_2= \begin{vmatrix}
 4 & 5 \\
 0& 6
\end{vmatrix}+\begin{vmatrix}
 1& 3\\
 0& 6
\end{vmatrix}+\begin{vmatrix}
 1& 2\\
 0& 4
\end{vmatrix}"


"=24+6+4=34"




"D_3=det\>|t|=1(24-0)+2(0-0)+3(0-0)"


"=24"




Characteristic polynomial:

"\lambda^3-11\lambda^2+34\lambda-24=(\lambda-1)(\lambda-4)(\lambda-6)"





The minimum polynomial will have roots

"1,4\>and\>6"


Putting the matrix in the characteristic polynomial


"T-1=\begin{pmatrix}
 0&2&3 \\
 0&3&5\\
0&0&5
\end{pmatrix}\>"



"T-4=\begin{pmatrix}
 -3& 2&3 \\
 0&0 & 5\\
0&0&2
\end{pmatrix}"




"T-6=\begin{pmatrix}
 -5&2& 3\\
 0&-2& 5\\
0&0&0
\end{pmatrix}"




"\begin{pmatrix}
 0&2 & 3\\
 0&3& 5\\
0&0&5
\end{pmatrix}\begin{pmatrix}
 -3& 2&3\\
 0&0&5\\
0&0&2
\end{pmatrix}\begin{pmatrix}
 -5&2&3 \\
 0&-2&5\\
0&0&0
\end{pmatrix}=\begin{pmatrix}
 0&0&0\\
 0&0& 0\\
0&0&0
\end{pmatrix}"



"\therefore" The minimal polynomial of "t" is


"M_tx=(x-1)(x-4)(x-6)"